# Survey of Hallucination in Natural Language Generation

### v6

@@ 13.4.1. Hallucination in Large Multimodal Models
Large multimodal models (LMMs) such as GPT-4V can generate rich and detailed response for visual inputs, and therefore being inherently riskier compared to previous VLMs for single-sentence captioning and short-answer VQA. Furthermore, images convey different levels of information encompassing object existence, attributes (color, shape, etc), spatial relationships, and sometimes high-level emotional responses they elicit (peaceful, beautiful, etc.), which makes mitigating hallucination in LMM even more challenging. The exploration of this area is still in its early stage, with a primary focus on object existence hallucination (Li et al., 2023c). However, it is imperative for LMMs to analyze images well beyond that surface level, otherwise the use of LMM would be redundant and a simple object detector would suffice. The most crucial factor in this problem is the availability and quality of data and supervision. Language-supervised representations (e.g., CLIP) have demonstrated inherent shortcomings in recognizing certain types of visual patterns, such as object orientation, quantity, and viewpoint (Tong et al., 2024). Even with perfect representation, imperfect vision-language alignment can result in a significant gap between the visual recognition abilities of the LMM and its visual backbone (Zhai et al., 2023). The cost of annotation during the mitigation of multimodal hallucination should also be considered. While there are works on the visual extension of faithfulness-oriented RLHF (Yu et al., 2023b), these essentially involve labeling more data, which is expensive and difficult to scale. The versatility of the visual domains also makes it challenging to guarantee robustness and generalization. Large-scale self-supervision may be the path forward, as supervised learning has proven to be insufficient for both the CV and NLP in the past decade.
@@ 13.4.2. Hallucination in Long-tail and Low-resource Domains
Hallucinations in LLMs in long-tail and low-resource domains still remains a significant challenge. This has become a new research focus – for example, many recent hallucination benchmarks are proposed to measure LLM hallucination in various long-tail and low-resource domains (see Section 13.2.1). However, our scope should not be confined to only narrow professional areas (e.g., medical, legal, finance), but should also consider broader scenarios which are also long-tailed and low-resource. For example, hallucination in low-resource languages is a less explored area, as hallucination mitigation methods applied to rich-resource languages do not necessarily generalize to low-resource ones. Similarly, the temporal axis also impacts the frequency of world knowledge appearance, where more recent data could also be considered as long-tailed and low-resource. Furthermore, beyond textual data, multi-modal scenarios are inherently low-resource for text-only pretrained LLMs. In light of the above, instead of developing domain-specific methods for hallucination mitigation, we need to design unified framework and generalizable strategies for addressing hallucination in various imbalanced scenarios.
@@ 13.4.5. Understanding Hallucination in LLMs
Although the research community has put extensive efforts to develop empirical methods for measuring and mitigating hallucinations, our understanding of LLM hallucination still remains limited. The absence of reliable theoretical frameworks and rigorous mathematical formulations hinders our ability to answer fundamental questions and make further advancements. For instance, how does hallucination correlate with model and data size (e.g., scaling laws)? What is the relationship between hallucination and auto-regressive generation? Does SFT and RLHF employ different mechanism to mitigate hallucinations? What is the lower-bound of hallucination across different tasks and domains? These questions are also intrinsically linked to a broader question – although now there are a lot of effort on empirically investigating what LLM can and cannot do from different dimensions, we still don’t understand why and how it works. There has been some attempts from various perspectives, such as knowledge encoding (Geva et al., 2021), compression (Delétang et al., 2023), model selection (Yadlowsky et al., 2023), compositionality (Dziri et al., 2023), computational perspective (Feng et al., 2023), but we are still a long way from truly understanding the mechanism behind various LLM behaviours.

### v7

@@ 13.4.1. Hallucination in Large Multimodal Models
Large multimodal models (LMMs) such as GPT-4V can generate rich and detailed responses for visual inputs, therefore being inherently riskier compared to previous VLMs for single-sentence captioning and short-answer VQA. Furthermore, images convey different levels of information encompassing object existence, attributes (color, shape, etc), spatial relationships, and sometimes high-level emotional responses they elicit (peaceful, beautiful, etc.), which makes mitigating hallucination in LMM even more challenging. The exploration of this area is still in its early stage, with a primary focus on object existence hallucination (Li et al., 2023c). However, it is imperative for LMMs to analyze images well beyond that surface level, otherwise, the use of LMM would be redundant and a simple object detector would suffice. The most crucial factor in this problem is the availability and quality of data and supervision. Language-supervised representations (e.g., CLIP) have demonstrated inherent shortcomings in recognizing certain types of visual patterns, such as object orientation, quantity, and viewpoint (Tong et al., 2024). Even with perfect representation, imperfect vision-language alignment can result in a significant gap between the visual recognition abilities of the LMM and its visual backbone (Zhai et al., 2023). The cost of annotation during the mitigation of multimodal hallucination should also be considered. While there are works on the visual extension of faithfulness-oriented RLHF (Yu et al., 2023b), these essentially involve labeling more data, which is expensive and difficult to scale. The versatility of the visual domains also makes it challenging to guarantee robustness and generalization. Large-scale self-supervision may be the path forward, as supervised learning has proven to be insufficient for both the CV and NLP in the past decade.
@@ 13.4.2. Hallucination in Long-tail and Low-resource Domains
Hallucinations in LLMs in long-tail and low-resource domains still remain a significant challenge. This has become a new research focus – for example, many recent hallucination benchmarks are proposed to measure LLM hallucination in various long-tail and low-resource domains (see Section 13.2.1). However, our scope should not be confined to only narrow professional areas (e.g., medical, legal, finance), but should also consider broader scenarios that are also long-tailed and low-resource. For example, hallucination in low-resource languages is a less explored area, as hallucination mitigation methods applied to rich-resource languages do not necessarily generalize to low-resource ones. Similarly, the temporal axis also impacts the frequency of world knowledge appearance, where more recent data could also be considered as long-tailed and low-resource. Furthermore, beyond textual data, multi-modal scenarios are inherently low-resource for text-only pretrained LLMs. In light of the above, instead of developing domain-specific methods for hallucination mitigation, we need to design a unified framework and generalizable strategies for addressing hallucination in various imbalanced scenarios.
@@ 13.4.5. Understanding Hallucination in LLMs
Although the research community has put extensive efforts into developing empirical methods for measuring and mitigating hallucinations, our understanding of LLM hallucination still remains limited. The absence of reliable theoretical frameworks and rigorous mathematical formulations hinders our ability to answer fundamental questions and make further advancements. For instance, how does hallucination correlate with model and data size (e.g., scaling laws)? What is the relationship between hallucination and auto-regressive generation? Do SFT and RLHF employ different mechanisms to mitigate hallucinations? What is the lower bound of hallucination across different tasks and domains? These questions are also intrinsically linked to a broader question – although now there are a lot of effort on empirically investigating what LLM can and cannot do from different dimensions, we still don’t understand why and how it works. There has been some attempts from various perspectives, such as knowledge encoding (Geva et al., 2021), compression (Delétang et al., 2023), model selection (Yadlowsky et al., 2023), compositionality (Dziri et al., 2023), computational perspective (Feng et al., 2023), but we are still a long way from truly understanding the mechanism behind various LLM behaviors.